OpenStudy (anonymous):

what is the excluded value? 3x(_over)/1-3x

OpenStudy (anonymous):

\[\left(\begin{matrix}3x \\ 1-3x\end{matrix}\right)\]

OpenStudy (anonymous):

x=1/3

OpenStudy (radar):

x=1/3 would be exc;ided as ot wwould result in division by zero

OpenStudy (anonymous):

1/3

OpenStudy (anonymous):

\[\frac{3x}{1-3x}\] is only valid when the denominator (bottom) is not zero. So the excluded value is when 1 - 3x = 0.

OpenStudy (anonymous):

check out the graph: http://bit.ly/kmZoIC You can see at x=1/3, it "blows up." It will approach negative infinity coming from one direction and positive infinity coming from the other. At exactly x=1/3, the function is undefined.

OpenStudy (anonymous):

so its 1over3?

OpenStudy (anonymous):

yes.

OpenStudy (radar):

Nice graph, I bookmarked that page.

OpenStudy (anonymous):

desmos.com is pretty awesome for a lot of questions here...